OpenStudy (anonymous):

point-slope form (7,13) (13,3)

OpenStudy (luigi0210):

Do you know the point-slope formula?

OpenStudy (anonymous):

point slope form for (x1,y1) (x2,y2) is y-y1 = m(x-x1) where m = (y2-y1)/(x2-x1)